OpenStudy (anonymous):

How many combinations of 5 people can be made from a group of 11 people if order doesn't matter?

OpenStudy (anonymous):

here are the answer choices: A) 39,916,800 B) 720 C) 462 D) 120

OpenStudy (anonymous):

help..

OpenStudy (anonymous):

@rayman5766

OpenStudy (rayman5766):

ok let me come back to this i will have some one help you ok

OpenStudy (anonymous):

ok but can you please hurry i need this answer

OpenStudy (rayman5766):

i will @sleepyhead314 can you help this kind man i am in a hurry my self

OpenStudy (anonymous):

are you talking to me or yourself?

OpenStudy (rayman5766):

well here is sleepy jead the dud that helped me out a lot of times maby he can help you

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (rayman5766):

ok so look he said he cant help but don't worry i have more people that can help you bud @texaschic101 pls help out this sir as i am trying to do my own work

OpenStudy (anonymous):

ok thanks

OpenStudy (rayman5766):

noprblem this is the smartest person i know so she should show up just give it a minute

OpenStudy (texaschic101):

sorry...I am not good at combinations or permutations

OpenStudy (anonymous):

it's ok. ill just guess

OpenStudy (anonymous):

and, future people... i feel sad for you because you'll have no answer!

OpenStudy (texaschic101):

5 people from a group of 11 (11 * 10 * 9 * 8 * 7) / (5 * 4 * 3 * 2 * 1) = 55440/120 = 462 http://betterexplained.com/articles/easy-permutations-and-combinations/